natalie us carrying six pages of nath homework and four pages of english homework. A gust of wind blows the pages out of her hands and she is only able to recover six random pages. What is the probability that she recovers all of her math homework.
Answer by Alan3354(69443)

natalie us carrying six pages of nath homework and four pages of english homework. A gust of wind blows the pages out of her hands and she is only able to recover six random pages. What is the probability that she recovers all of her math homework.
-----------
A total of 10 pages.
The 1st is 6/10
Then 5/9, 4/8, 3/7, 2/6, 1/5
-----
It's the product:
= 6/10, 5/9, 4/8, 3/7, 2/6, 1/5
= 3/5, 5/9, 1/2, 3/7, 1/3, 1/5
= 1/210